# Do It Yourself: Learning Semantic Correspondence from Pseudo-Labels

### v2

@@ 4.1 Experimental Results
The category-specific results in Tab. 2 indicate that our model improves significantly for almost all categories compared to the current SOTA, showcasing that our pseudo-labeling strategy is effective, resulting in a new SOTA for the SPair-71k dataset. We observe particularly large improvement for categories that are symmetric or contain repeated object parts, e.g., $+15.7\%$ for bus or $+14.0\%$ for car. While Mariotti et al. [35] already significantly improved the performance for such categories compared to a naïve zero-shot strategy, we further outperform their approach. This is, on the one hand, thanks to our proposed training strategy using pseudo-labels, which effectively learns to combine the advantages of the complementary SD and DINOv2 features. Secondly, in contrast to Mariotti et al. [35], we do not modulate the original features but we we only reject matches that are far apart when mapped to a spherical prototype.
@@
Additionally, we also evaluate on the AP-10k dataset and report the results in Tab. 3. We find that, despite not being trained on this dataset and only having a few animal categories in the SPair-71k training dataset, we outperform the current SOTA, the geometry-aware approach by Zhang et al. [63], showcasing the generalization capabilities of our method. We additionally compare to a fully-supervised model [63] trained on SPair-71k. We observe a severe performance drop of this model ($-17.4\%$ PCK@0.1 for I.S.) compared to fully-supervised training on AP-10k [63, Tab.3]. On the other hand, our approach does not suffer such a performance drop and, thus, achieves comparable results to a model trained with manual annotations. We hypothesize that this is due to the denser and more diverse pseudo-labels compared to the fixed set of manual annotated keypoints that are being overfitted in supervised training.
@@
Our weakly-supervised training strategy allows us to scale the training to larger object-centric datasets that do not include keypoint annotations. While pre-training improves the model performance (Tab. 3 and Tab. 4) on SPair-71k and AP-10k, we surprisingly even outperform the previous SOTA on SPair-71k when not trained on the SPair-71k dataset, showcasing strong generalization capabilities. This is in contrast to the naïve weighted application of the ImageNet-3D spherical mapper, where the results significantly deteriorate, potentially explained by the fact that this model does not generalize well to the unseen categories in SPair-71k. We also study the impact of automatically acquired 3D pose labels in the supplementary Sec. A.3.
@@
We present qualitative results in Fig. 4 comparing our model to current SOTA models on four challenging examples of the SPair-71k test dataset. While the spherical mapper effectively addresses cases of symmetries and repeated object parts for rigid objects, it comes with the inherent challenge of finding the correct weighting factor for interpolating between foundational and processed spherical feature similarities where the optimal weight depends on the considered category and sample. For example, while the weighting is too low in (a), the weighting is too large in (d) since the ill-defined background features of SphMap result in a match with an object in the background. Additionally, its effectiveness is limited to categories that are well represented by a spherical prototype, e.g., a potential reason for the motorbike failure case in (c). TLR and DistillDIFT, both, suffer from the fact that simply flipping an object does not resolve all ambiguities, such as for categories with repeated object parts (a,b), rotated objects (b), or cases with left-/right ambiguities (d). DistillDIFT is less error-prone to wrong matches caused by background clutter (c), where TLR and SphMap suffer from the fact that they rely on the potentially noisy nearest-neighbor of SD and DINOv2 features. However, DistillDIFT still fails to disambiguate the human and the horse in (d) and it suffers from left-right ambiguity if the flipping is not effective (d). Furthermore, as visible in (b), DistillDIFT’s performance deteriorates for pairs with heavy appearance change, which can be possibly attributed to the multi-view fine-tuning on 3D instances, which does not generalize well to cross-instance correspondences. In contrast, our method finds high-quality semantic correspondences for these challenging examples since it successfully resolves feature ambiguities and does not rely on the weighting of two similarity terms.
@@ 4.2 Ablations
Failure Cases. While our method outperforms previous methods, it still fails for some challenging cases of SPair-71K, such as images containing objects that are upside-down. We show visual examples and discuss failure cases of our method in Sec. A.2.
@@ 5 Conclusion
Limitations and future work. Although our weak supervision strategy scales to larger datasets, reducing the supervision requirements is of importance to further scale the approach. See Sec. A.3 for a preliminary feasibility analysis. Furthermore, while the match rejection with the spherical mapper shows significant improvements, it is less effective for objects that are not represented well by a sphere, which motivates extending this strategy. We hope that the proposed improved correspondence features can serve to improve tasks that rely on semantic features.
Acknowledgements. Adam Kortylewski acknowledges support via his Emmy Noether Research Group funded by the German Research Foundation (DFG) under Grant No. 468670075. Thomas Wimmer is supported through the Max Planck ETH Center for Learning Systems.
Appendix A Appendix
A.1 Details on ImageNet-3D Trained Model
@@
For training the spherical mapper, we use the same hyperparameter as [35] and we train for 200 epochs on the ImageNet-3D dataset. The generation of the pseudo-labels and training of the adapter follow the same hyperparameters as our presented model for SPair-71k. We train on this larger dataset for 400k steps.
A.2 Discussion of Failure Cases
@@
A.3 Training with Less Supervision
@@
A.4 Pre-Training with Pseudo-Labels
@@
A.5 Pseudo-Label Generation without SD
@@
A.6 Ablation of Number of Feature Channels
Learning a refining module allows reducing the number of channels of the features used for nearest neighbor computation. The performance does not heavily drop, which might be a valuable option for memory-constrained applications.
@@ A.6 Ablation of Number of Feature Channels
A.7 More Qualitative Results

### v3

@@ 4.1 Experimental Results
The category-specific results in Tab. 2 show that our model achieves strong gains across nearly all categories, setting a new SOTA on SPair-71k and demonstrating the effectiveness of our pseudo-labeling strategy. Improvements are especially large for categories that are symmetric or contain repeated object parts, e.g., $15.7\%$ for bus or $14.0\%$ for car. While Mariotti et al. [35] already improved performance for such cases, our approach outperforms theirs by effectively combining SD and DINOv2 features through pseudo-label supervision. Unlike their method, we retain original features and simply reject matches that strongly deviate from a spherical prototype.
@@
Additionally, we also evaluate on the AP-10k dataset and report the results in Tab. 3. We find that, despite not being trained on this dataset and only having a few animal categories in the SPair-71k training dataset, we outperform the current SOTA, the geometry-aware approach by Zhang et al. [63], showcasing the generalization capabilities of our method. We additionally compare to a fully-supervised model [63] trained on SPair-71k. We observe a severe performance drop of this model ($17.4\%$ PCK@0.1 for I.S.) compared to fully-supervised training on AP-10k [63, Tab.3]. On the other hand, our approach does not suffer such a performance drop and, thus, achieves comparable results to a model trained with manual annotations. We hypothesize that this is due to the denser and more diverse pseudo-labels compared to the fixed set of manually annotated keypoints that are being overfitted in supervised training.
@@
Our weakly-supervised training strategy allows scaling the training to larger object-centric datasets that do not include keypoint annotations. While pre-training improves the model performance on SPair-71k and AP-10k (Tab. 3 and Tab. 4), we surprisingly even outperform the previous SOTA on SPair-71k despite not having trained on the SPair-71k dataset, showcasing strong generalization capabilities. This is in contrast to the naïve weighted application of the ImageNet-3D spherical mapper, where the results significantly deteriorate, potentially explained by the fact that this model does not generalize well to the unseen categories in SPair-71k. We also study the impact of automatically acquired 3D pose labels in the supplementary.
@@
We present qualitative results in Fig. 4 comparing our model to current SOTA models on four challenging examples of the SPair-71k test dataset. While the spherical mapper effectively addresses cases of symmetries and repeated object parts for rigid objects, it comes with the inherent challenge of finding the correct weighting factor for interpolating between foundational and processed spherical feature similarities, where the optimal weight depends on the considered category and sample. For example, while the weighting is too low in (a), the weighting is too large in (d) since the ill-defined background features of SphMap result in a match with an object in the background. Additionally, its effectiveness is limited to categories that are well represented by a spherical prototype, e.g., a potential reason for the motorbike failure case in (c). TLR and DistillDIFT both suffer from the fact that simply flipping an object does not resolve all ambiguities, such as for categories with repeated object parts (a,b), rotated objects (b), or cases with left-/right ambiguities (d). DistillDIFT is less prone to incorrect matches caused by background clutter (c), whereas TLR and SphMap suffer from the fact that they rely on the potentially noisy nearest neighbor of SD and DINOv2 features. However, DistillDIFT still fails to disambiguate the human and the horse in (d), and it suffers from left-right ambiguity if the flipping is not effective (d). Furthermore, as visible in (b), DistillDIFT’s performance deteriorates for pairs with heavy appearance change, which may be attributed to the multi-view fine-tuning on 3D instances, which does not generalize well to cross-instance correspondences. In contrast, our method identifies high-quality semantic correspondences for these challenging examples, as it successfully resolves feature ambiguities and does not rely on the weighting of two similarity terms.
@@ 4.2 Ablations
Failure Cases.
While our method outperforms previous methods, it still fails for some challenging cases of SPair-71K, such as images containing objects that are upside-down. We show visual examples and discuss failure cases of our method in the supplementary.
@@ 5 Conclusion
Limitations and future work.
Although our weak supervision strategy scales to larger datasets, reducing the supervision requirements is of importance to further scale the approach. See the supplementary for a preliminary feasibility analysis. Furthermore, while the match rejection with the spherical mapper shows significant improvements, it is less effective for objects that are not represented well by a sphere, which motivates extending this strategy. We hope that the proposed improved correspondence features can enhance tasks that rely on semantic features.
Acknowledgments
Adam Kortylewski acknowledges support via his Emmy Noether Research Group funded by the German Research Foundation (DFG) under Grant No. 468670075. Thomas Wimmer is supported through the Max Planck ETH Center for Learning Systems.
Supplementary Material
Appendix A Details on ImageNet-3D Trained Model
@@
For training the spherical mapper, we use the same hyperparameter as Mariotti et al. [35] and we train for 200 epochs on the ImageNet-3D dataset. The generation of the pseudo-labels and training of the adapter follow the same hyperparameters as our presented model for SPair-71k. We train on this larger dataset for 400k steps.
Appendix B Discussion of Failure Cases
@@
Appendix C Training with Less Supervision
@@
Appendix D Pre-Training with Pseudo-Labels
@@
Appendix E Pseudo-Label Generation without SD
@@
Appendix F Ablation of Number of Feature Channels
Learning a refining module allows reducing the number of channels of the features used for nearest neighbor computation. The performance does not heavily drop with a lower number of channels (Tab. 8), which might be a valuable option for memory-constrained applications.
Appendix G More Qualitative Results